# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ XII-C A unified FA framework
FL operates under a unified framework to embed similar computation procedures and insight structures when training various types of neural networks. Such a unified framework plays a vital role in boosting the wide studies and applications of FL. On the other hand, FA mechanisms designed so far are still highly task-specific. While it seems intuitive that different data analytics tasks are naturally different regarding their computation procedures, data structures, etc, there already exists some efforts to provide a unified framework for a particular class of data science problem[16]. Attracted by the potential benefits of a unified framework, it remains a grand open problem to design a unified framework for broader classes of data analytics tasks or even universal data analytics tasks. Specifically, motivated by this, some open questions and interesting directions could be (i) What are the proper architecture, and interface design for a unified FA framework? (ii) If we still adopt the current local insight upload and global aggregation scheme, how to design a general aggregator or general insight form that capable of incorporating diverse requirements? (iii) Witnessing the advancement in generative models, is it possible to use generative models at the server side, essentially, as a general aggregator, to further reduce the reliance on locally uploaded insights so that the communication cost can be reduced further and versatile privacy-preserving tasks can be supported.
XII-D System-efficient FA
@@ XII-E Wireless communication for FA
Over-the-air computation (AirComp) leverages the superposition property of wireless channels to enable simultaneous transmission and aggregation of signals from multiple devices [68]. It has been applied to FL by aggregating model updates from numerous devices directly over the air. It not only enhances communication efficiency but also maintains the privacy of clients’ local data, addressing key bottlenecks in traditional FL approaches, such as limited bandwidth and latency issues. How FA could benefit from physical layer techniques also presents interesting and open challenges.
XIII Conclusion
@@
Due to the exponential growth of edge data and the growing awareness of data privacy, privacy-preserving distributed data processing has attracted wide interests from both academia and industry. Federated analytics is an emerging collaborative data processing framework for descriptive data science tasks without centralizing the raw data. It brings significant benefits in privacy protection, communication reduction, and task coverage. Although FA has been widely studied in industry and academia, a systematic review of the existing efforts in FA has not been conducted yet. This survey fills the gap by first comprehensively reviewing the key concepts in FA and its relationship with similar techniques. It then presents a detailed taxonomy to categorize FA studies from both applications and system characteristics. Key challenges and enabling techniques are introduced with a specific focus on privatization, analytics, and optimization. A wide spectrum of FA tasks are then reviewed demonstrating the generality of FA applications. Finally, we discuss the open issues and future directions in FA, from the perspective of application, privacy protection, framework, system optimization, and cross-layer design. This survey summarizes the existing efforts in FA and the huge intersection among FA and data science, privacy, distributed computing, wireless communications, and networking systems. Overall, FA research is widely interdisciplinary, approachable, and contains many critical problems that would benefit greatly from the expertise of all related areas.

### v3

@@ XI-B A unified FA framework
FL operates under a unified framework to embed similar computation procedures and insight structures when training various types of neural networks. Such a unified framework plays a vital role in boosting the wide studies and applications of FL. On the other hand, FA mechanisms designed so far are still highly task-specific. While it seems intuitive that different data analytics tasks are naturally different regarding their computation procedures, data structures, etc, there already exists some efforts to provide a unified framework for a particular class of data science problem[26]. Some FA solutions for database operations Attracted by the potential benefits of a unified framework, it remains a grand open problem to design a unified framework for broader classes of data analytics tasks or even universal data analytics tasks. Specifically, motivated by this, some open questions and interesting directions could be (i) What are the proper architecture, and interface design for a unified FA framework? (ii) If we still adopt the current local insight upload and global aggregation scheme, how to design a general aggregator or general insight form that capable of incorporating diverse requirements? (iii) Witnessing the advancement in generative models, is it possible to use generative models at the server side, essentially, as a general aggregator, to further reduce the reliance on locally uploaded insights so that the communication cost can be reduced further and versatile privacy-preserving tasks can be supported.
XI-C Privacy preservation at scale
As FA systems grow in scale with more data or users involved, ensuring privacy becomes more challenging. This challenge is exacerbated when analysts seek high-accuracy analytics. Existing privacy-preserving techniques exhibit distinct characteristics concerning privacy assurance, data utility, and scalability. For example, DP has long been criticized for offering a poor accuracy-privacy tradeoff, prohibiting its applications in domains that request accurate population-level profiling or analytics. Other cryptographic techniques, such as MPC, encounter difficulties in scaling to practical sizes for mobile applications. Consequently, FA mechanisms have to be carefully designed to strike a delicate balance between privacy, utility, and scalability. Some recent concepts regarding privacy preservation, like private cloud computing, are promising to enhance FA algorithmic design: it might be possible to decompose the FA algorithm into local computation part and private cloud computation part, where the local computation part conducts lightweight computations that extract insights to reduce communication overheads, and the computation-intensive schemes, such as cryptography-based privacy preservation, is conducted on the private cloud.
XI-D Formal measurement of privacy preservation
Existing FA works utilize task-specific algorithms, and provide different privacy guarantees. These guarantees are under different threat models and assumptions and eventually provide different privacy preservation that are hard to trace and compare. Taking DP as an example, in DP-based FA solutions, in addition to conventional CDP/LDP separation, various DP definitions (e.g., sequence DP, token DP) are designed with different focuses on privacy preservation. As a result, even for two FA schemes both providing DP guarantees with the same values of $\epsilon$ and $\delta$, their strength in privacy preservation may vary a lot. Therefore, it is crucial to provide a formal and comprehensive measure of privacy preservation, covering both DP, cryptography, and other privatization approaches. Such a formal measurement is beneficial for the development of FA platforms where FA are provided as a service, and also helpful for the ambition of a unified FA framework. Possible approaches include quantifying privacy preservation based on the performance of adversarial privacy attacks. However, such an approach lacks theoretical justification and is limited by attack methods and models.
XI-E System-efficient FA
@@ XI-F Wireless communication for FA
Over-the-air computation (AirComp) leverages the superposition property of wireless channels to enable simultaneous transmission and aggregation of signals from multiple devices [112, 113, 199, 200]. It has been applied to FL by aggregating model updates from numerous devices directly over the air. It not only enhances communication efficiency but also maintains the privacy of clients’ local data, addressing key bottlenecks in traditional FL approaches, such as limited bandwidth and latency issues. However, the existing research on AirComp focuses on FL rather than FA. Since the model aggregation in FL usually are in a simpler form (FedAvg and its variations), FA aggregation usually requires a more complex computation procedure. How the idea of AirComp can be applied for complex aggregation schemes, and how FA could benefit from physical layer techniques, also present interesting and open challenges.
XI-G Resource management in FA
Essentially as an edge computing scheme, FA should naturally applied to all kinds of resource-limited scenarios. The deployment of FA is companied by all kinds of resource constraints: the local computation and insight aggregation phases require computation resources; the model and insight transmission require communication resources; the level of client data leakage consumes the novel privacy resource at the client side; even the participation of client can also be considered as a kind of resource to some extent. Significant research efforts have been made in resource management and resource saving in FL, including all these aspects of computation [201], communication [85], privacy [114], and client participation [202]. However, designing novel mechanisms for resource management and saving for FA, or adapting these FL mechanisms into FA applications, is a non-trivial unexplored topic. Many existing schemes for FL may highly rely on the FL architecture and are not applicable in FA applications. For example, gradient compression and quantization are widely used in communication-efficient FL but are unlikely to be useful in FA since the uploaded insights are not in the form of gradient. Therefore, designing resource management and saving schemes for FA is a critical issue for its wide deployment.
XII Conclusion
@@
Due to the exponential growth of edge data and the growing awareness of data privacy, privacy-preserving distributed data processing has attracted wide interest from both academia and industry. FA is an emerging collaborative data processing framework for descriptive data science tasks without centralizing the raw data. It brings significant benefits in privacy protection, communication reduction, and task coverage. Although FA has been widely studied in industry and academia, a systematic review of the existing efforts in FA has not been conducted yet. This survey fills the gap by first comprehensively reviewing the key concepts in FA, its relationship with similar techniques, and its key challenges. It presents a detailed taxonomy to categorize FA studies from both applications and system characteristics. Key enabling techniques are introduced with a specific focus on privatization, analytics, and optimization. A wide spectrum of FA tasks are then reviewed demonstrating the generality of FA applications. Finally, we discuss the open issues and future directions in FA, from the perspective of application, privacy protection, framework, system optimization, and cross-layer design. This survey summarizes the existing efforts in FA and the huge intersection among FA and data science, privacy, distributed computing, wireless communications, and networking systems, and derives key lessons learned that represent our insights on this field. Overall, FA research is widely interdisciplinary, approachable, and contains many critical problems that would benefit greatly from the expertise of all related areas.